# MuSciClaims: Multimodal Scientific Claim Verification

### v1

@@ 7 Conclusion
Assessing whether claims are supported requires understanding the methods and data presented in associated figures.
One must find the correct piece of information in the figure and then combine it with the caption.
This paper introduces
MuSciClaims, a new diagnostic dataset to evaluate the claim verification capabilities of VLMs.
We find that most VLMs are poor at this task out-of-the-box, and chain-of-thought only helps slightly.
Particularly, they are significantly worse at understanding that given evidence contradicts (or is not related to) the claim.
EvidenceLocalization shows that models are bad at identifying the right panel of data, a critical flaw in their claim verification capabilities.
Cross-ModalAggregation indicates that models do not effectively use both visual and textual information for their judgments.
In fact, BasicVisualUnderstanding reveals that they do not understand some obvious characteristics of the associated figures.
Our results establish the current abilities of VLMs for claim verification over heterogeneous, information-rich scientific figures, and our diagnostics highlight specific avenues of research to improve them.
@@
In this work, we do not evaluate the reasoning produced by VLMs. Such evaluation requires experts with incredibly specific domain expertise. Even a graduate student (PhD level) or faculty cannot verify reasoning for all domains covered in Cell. For instance, an expert in ecology cannot easily judge the reasoning about claims in cellular biology. We will explore how to conduct better evaluations as part of future work.
While our work only considers scientific papers in the life sciences, our methods can in principle be applied to many other domains. Physical sciences such as physics and chemistry, among others, are domains worth investigating. Our work only investigates English-language documents and this limits the generalizability of our findings to other languages, although most scientific articles are disseminated in English.
Due to high cost of the recently released o4 models, we are unable to analyze it across the full spectrum of our diagnostics.
@@ Acknowledgments
This material is based on research that is in part supported by the DARPA for the SciFy program under agreement number HR00112520301. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either express or implied, of DARPA or the U.S. Government.
@@ Appendix C Automatic Extraction Details
To construct a high-quality multimodal benchmark for scientific claim verification, we developed an automated pipeline for extracting textual claims and their associated visual elements from research articles. Our approach operates over full-text HTML and PDF documents sourced from Cell (cell.com), a leading biomedical journal series. This creates reliable mapping between complex scientific assertions and their supporting visual evidence with minimal manual supervision. Figure 7 shows an overview of our dataset creation process. During dataset collection, no personal identifying info (PII) was collected. None of the collected data contained any offensive content.
MuSciClaims
@@ C.1 Automatic Figure Extraction
We extract figures and their corresponding captions from the structured HTML versions of articles sourced from cell.com. Each article contains embedded figure blocks that follow consistent filename conventions and DOM structures, allowing for reliable identification and extraction. Figures are mapped to canonical identifiers (e.g., figure_1, figure_2, etc.) to ensure consistency across the dataset.
@@ Appendix D Model Setup
Python was the main scripting language for data collection and experimentation. For experiments using closed source models, we used OpenAI555https://openai.com/api/pricing/ and Anthropic666https://://www.anthropic.com/pricing APIs. The total cost for OpenAI was 270 USD and 150 USD for Claude. The open-source experiments were conducted on A6000 GPUs, each having GB. The total GPU hours for all the experiments was 28. The models were downloaded from Huggingface and hosted for inference using Huggingface transformers module and vLLM. We use GitHub Co-Pilot to help with writing code but verify it manually before running any experiments.
Appendix E Additional Analysis Results
| Support | NonSupport | MuSciClaims | ||||||||
@@ Appendix E Additional Analysis Results
MuSciClaims task as a two-class problem
Table 6 presents the results when the main claim verification task of
Panel Complexity
Appendix F Prompts Used

### v2

@@ 7 Conclusion
Assessing whether claims are supported requires understanding the methods and data presented in associated figures. One must find the correct piece of information in the figure and then combine it with the caption. This paper introduces MuSciClaims, a new diagnostic dataset to evaluate the claim verification capabilities of VLMs. We find that most VLMs are poor at this task out-of-the-box, and chain-of-thought only helps slightly. Particularly, they are significantly worse at understanding that given evidence contradicts (or is not related to) the claim. EvidenceLocalization shows that models are bad at identifying the right panel of data, a critical flaw in their claim verification capabilities. Cross-ModalAggregation indicates that models do not effectively use both visual and textual information for their judgments. Diagnostics also reveal that they do not understand some obvious characteristics of the associated figures. Our results establish the current abilities of VLMs for claim verification over heterogeneous, information-rich scientific figures, and our diagnostics highlight specific avenues of research to improve them.
@@
Due to the difficulty of creating data that VLMs have not already seen (published after their cutoff dates), we are unable to train models for this task. We perform few-shot experiments with closed-source models (4o-mini, 4o and Sonnet) but we leave further exploration of different methods of example selection to future work.
@@
In this work, we perform some qualitative evaluation of weaknesses in the reasoning produced by VLMs. However, we are unable to do so at scale. Such evaluation requires experts with incredibly specific domain expertise. Even a graduate student (PhD level) or faculty cannot verify reasoning for all domains covered in Nature Physics, Journal of the American Chemical Society and Cell. For instance, an expert in ecology cannot easily judge the reasoning about claims in cellular biology. We will explore how to conduct better evaluations as part of future work.
Our work only investigates English-language documents and this limits the generalizability of our findings to other languages, although most scientific articles are disseminated in English.
Due to high cost of the recently released o3 and o4-mini models, we are unable to analyze it across the full spectrum of our diagnostics.
@@ Acknowledgments
This material is based on research that is in part supported by the DARPA for the SciFy program under agreement number HR00112520301. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either express or implied, of DARPA or the U.S. Government. The authors would also like to thank Aakanksha Rajiv Kapoor for her help with understanding the structure and content of papers published in the Cell journal and for help with the qualitative error analysis of model outputs.
@@ Appendix A Benchmark Models
o3-2025-04-16
excels at solving complex math, coding, and scientific challenges while demonstrating strong visual perception and analysis. It uses tools in its chains of thought to augment its capabilities; for example, cropping or transforming images, searching the web, or using Python to analyze data during the thought process.
@@ Appendix C Automatic Extraction Details
To construct a high-quality multimodal benchmark for scientific claim verification, we developed an automated pipeline for extracting textual claims and their associated visual elements from research articles. Our approach operates over full-text HTML and PDF documents sourced from Nature Physics (https://www.nature.com/nphys/), the Journal of the American Chemical Society (https://pubs.acs.org/journal/jacsat) and Cell (https://www.cell.com), leading impact factor venues in their respective fields. It creates reliable mapping between complex scientific assertions and their supporting visual evidence with minimal manual supervision. During dataset collection, no personal identifying info (PII) was collected. None of the collected data contained any offensive content.
@@ C.1 Automatic Figure Extraction
We extract figures and their corresponding captions from the structured HTML versions of articles sourced from these journals. Each article contains embedded figure blocks that follow consistent filename conventions and DOM structures, allowing for reliable identification and extraction. Figures are mapped to canonical identifiers (e.g., figure_1, figure_2, etc.) to ensure consistency across the dataset.
@@ Appendix D Model Setup
Python was the main scripting language for data collection and experimentation. For experiments using closed source models, we used OpenAI555https://openai.com/api/pricing/ and Anthropic666https://://www.anthropic.com/pricing APIs. The total cost for OpenAI was 400 USD and 160 USD for Claude. The open-source experiments were conducted on A6000 GPUs, each having GB. The total GPU hours for all the experiments was 40. The models were downloaded from Huggingface and hosted for inference using Huggingface transformers module and vLLM. We use GitHub Co-Pilot to help with writing code but verify it manually before running any experiments.
Appendix E Additional Results
E.1 Model Performance by Domain
| Physics | Chemistry | Biology | Overall | ||||||||||
| P | R | F | P | R | F | P | R | F | P | R | F | ||
| 4o-mini | D | 0.56 | 0.45 | 0.38 | 0.65 | 0.55 | 0.49 | 0.60 | 0.47 | 0.42 | 0.60 | 0.48 | 0.43 |
| RD | 0.52 | 0.47 | 0.43 | 0.57 | 0.54 | 0.51 | 0.57 | 0.50 | 0.46 | 0.56 | 0.50 | 0.47 | |
| 4o | D | 0.53 | 0.47 | 0.43 | 0.69 | 0.59 | 0.57 | 0.73 | 0.54 | 0.52 | 0.68 | 0.54 | 0.51 |
| RD | 0.55 | 0.50 | 0.48 | 0.64 | 0.60 | 0.57 | 0.69 | 0.59 | 0.56 | 0.65 | 0.57 | 0.55 | |
| Sonnet | D | 0.62 | 0.53 | 0.50 | 0.68 | 0.62 | 0.61 | 0.76 | 0.68 | 0.68 | 0.71 | 0.65 | 0.64 |
| RD | 0.62 | 0.53 | 0.51 | 0.72 | 0.66 | 0.66 | 0.78 | 0.70 | 0.70 | 0.73 | 0.66 | 0.66 | |
| o3 | RD | 0.58 | 0.55 | 0.54 | 0.72 | 0.71 | 0.71 | 0.79 | 0.77 | 0.77 | 0.73 | 0.72 | 0.72 |
| o4-mini | RD | 0.64 | 0.59 | 0.57 | 0.73 | 0.72 | 0.71 | 0.81 | 0.77 | 0.77 | 0.75 | 0.72 | 0.72 |
| Phi-4 | D | 0.43 | 0.40 | 0.38 | 0.60 | 0.48 | 0.47 | 0.63 | 0.47 | 0.43 | 0.54 | 0.46 | 0.43 |
| RD | 0.62 | 0.37 | 0.27 | 0.52 | 0.39 | 0.32 | 0.60 | 0.42 | 0.37 | 0.58 | 0.41 | 0.34 | |
| Llava-Next | D | 0.28 | 0.39 | 0.30 | 0.36 | 0.46 | 0.37 | 0.74 | 0.42 | 0.32 | 0.68 | 0.42 | 0.33 |
| RD | 0.52 | 0.39 | 0.34 | 0.55 | 0.45 | 0.40 | 0.50 | 0.44 | 0.38 | 0.51 | 0.43 | 0.38 | |
| Llama-3.2 | D | 0.51 | 0.45 | 0.41 | 0.56 | 0.48 | 0.44 | 0.60 | 0.50 | 0.47 | 0.57 | 0.49 | 0.45 |
| RD | 0.53 | 0.41 | 0.35 | 0.52 | 0.41 | 0.34 | 0.60 | 0.43 | 0.36 | 0.56 | 0.42 | 0.35 | |
| Molmo | D | 0.46 | 0.41 | 0.37 | 0.55 | 0.47 | 0.42 | 0.62 | 0.50 | 0.46 | 0.57 | 0.47 | 0.43 |
| RD | 0.43 | 0.40 | 0.38 | 0.44 | 0.41 | 0.37 | 0.47 | 0.43 | 0.41 | 0.46 | 0.42 | 0.39 | |
| InternVL3 | D | 0.59 | 0.58 | 0.57 | 0.73 | 0.72 | 0.72 | 0.77 | 0.74 | 0.74 | 0.72 | 0.70 | 0.70 |
| RD | 0.56 | 0.49 | 0.46 | 0.69 | 0.57 | 0.56 | 0.75 | 0.59 | 0.58 | 0.69 | 0.57 | 0.55 | |
| Qwen2.5 | D | 0.68 | 0.58 | 0.54 | 0.70 | 0.65 | 0.62 | 0.72 | 0.66 | 0.65 | 0.70 | 0.65 | 0.63 |
| RD | 0.63 | 0.47 | 0.43 | 0.70 | 0.53 | 0.51 | 0.71 | 0.54 | 0.52 | 0.68 | 0.53 | 0.50 | |
| DeepSeek | D | 0.47 | 0.47 | 0.46 | 0.52 | 0.50 | 0.49 | 0.51 | 0.51 | 0.50 | 0.50 | 0.67 | 0.58 |
| RD | 0.42 | 0.42 | 0.39 | 0.46 | 0.46 | 0.42 | 0.50 | 0.47 | 0.46 | 0.48 | 0.46 | 0.44 | |
Table 6 presents model performance for different domains in MuSciClaims. On average, we note that models are worst at verifying Physics claims and best at judging Chemistry claims. However, the highest performance is achieved by o3 and o4-mini on Biology claims. We also present per-label metrics for model performance in each domain in Table 7, Table 8 and Table 9.
| Support | Neutral | Contradict | Overall | ||||||||||
| P | R | F | P | R | F | P | R | F | P | R | F | ||
| 4o-mini | D | 0.40 | 0.81 | 0.54 | 0.52 | 0.47 | 0.49 | 0.75 | 0.06 | 0.12 | 0.56 | 0.45 | 0.38 |
| RD | 0.41 | 0.79 | 0.54 | 0.56 | 0.46 | 0.50 | 0.58 | 0.15 | 0.23 | 0.52 | 0.47 | 0.43 | |
| 4o | D | 0.40 | 0.83 | 0.54 | 0.63 | 0.38 | 0.47 | 0.56 | 0.19 | 0.28 | 0.53 | 0.47 | 0.43 |
| RD | 0.44 | 0.71 | 0.54 | 0.54 | 0.57 | 0.56 | 0.66 | 0.22 | 0.33 | 0.55 | 0.50 | 0.48 | |
| Sonnet | D | 0.44 | 0.77 | 0.56 | 0.60 | 0.58 | 0.59 | 0.81 | 0.23 | 0.36 | 0.62 | 0.53 | 0.50 |
| RD | 0.45 | 0.79 | 0.57 | 0.60 | 0.54 | 0.57 | 0.81 | 0.26 | 0.39 | 0.62 | 0.53 | 0.51 | |
| o3 | RD | 0.59 | 0.52 | 0.55 | 0.50 | 0.79 | 0.62 | 0.63 | 0.34 | 0.45 | 0.58 | 0.55 | 0.54 |
| o4-mini | RD | 0.53 | 0.66 | 0.59 | 0.56 | 0.76 | 0.65 | 0.82 | 0.34 | 0.49 | 0.64 | 0.59 | 0.57 |
| Phi-4 | D | 0.37 | 0.70 | 0.49 | 0.43 | 0.24 | 0.31 | 0.48 | 0.27 | 0.35 | 0.43 | 0.40 | 0.38 |
| RD | 0.35 | 0.90 | 0.50 | 1.00 | 0.02 | 0.04 | 0.51 | 0.19 | 0.27 | 0.62 | 0.37 | 0.27 | |
| Llava-Next | D | 0.36 | 0.78 | 0.49 | 0.47 | 0.39 | 0.42 | 0.00 | 0.00 | 0.00 | 0.28 | 0.39 | 0.30 |
| RD | 0.37 | 0.74 | 0.49 | 0.40 | 0.35 | 0.38 | 0.80 | 0.08 | 0.15 | 0.52 | 0.39 | 0.34 | |
| Llama-3.2 | D | 0.40 | 0.80 | 0.53 | 0.54 | 0.41 | 0.46 | 0.59 | 0.14 | 0.22 | 0.51 | 0.45 | 0.41 |
| RD | 0.36 | 0.90 | 0.52 | 0.59 | 0.18 | 0.27 | 0.65 | 0.16 | 0.25 | 0.53 | 0.41 | 0.35 | |
| Molmo | D | 0.38 | 0.83 | 0.52 | 0.51 | 0.22 | 0.31 | 0.50 | 0.19 | 0.27 | 0.46 | 0.41 | 0.37 |
| RD | 0.37 | 0.72 | 0.49 | 0.50 | 0.23 | 0.31 | 0.43 | 0.26 | 0.32 | 0.43 | 0.40 | 0.38 | |
| InternVL3 | D | 0.55 | 0.69 | 0.61 | 0.59 | 0.68 | 0.63 | 0.63 | 0.38 | 0.47 | 0.59 | 0.58 | 0.57 |
| RD | 0.43 | 0.78 | 0.55 | 0.55 | 0.49 | 0.52 | 0.70 | 0.20 | 0.31 | 0.56 | 0.49 | 0.46 | |
| Qwen2.5 | D | 0.52 | 0.67 | 0.59 | 0.56 | 0.82 | 0.66 | 0.96 | 0.24 | 0.38 | 0.68 | 0.58 | 0.54 |
| RD | 0.40 | 0.84 | 0.54 | 0.55 | 0.39 | 0.45 | 0.94 | 0.18 | 0.30 | 0.63 | 0.47 | 0.43 | |
| DeepSeek | D | 0.52 | 0.43 | 0.47 | 0.48 | 0.72 | 0.57 | 0.41 | 0.27 | 0.33 | 0.47 | 0.47 | 0.46 |
| RD | 0.40 | 0.60 | 0.48 | 0.45 | 0.50 | 0.47 | 0.43 | 0.16 | 0.23 | 0.42 | 0.42 | 0.39 | |
| Support | Neutral | Contradict | Overall | ||||||||||
| P | R | F | P | R | F | P | R | F | P | R | F | ||
| 4o-mini | D | 0.46 | 0.86 | 0.60 | 0.70 | 0.69 | 0.69 | 0.79 | 0.11 | 0.19 | 0.65 | 0.55 | 0.49 |
| RD | 0.46 | 0.74 | 0.57 | 0.63 | 0.67 | 0.65 | 0.63 | 0.21 | 0.32 | 0.57 | 0.54 | 0.51 | |
| 4o | D | 0.46 | 0.91 | 0.61 | 0.91 | 0.59 | 0.72 | 0.70 | 0.25 | 0.37 | 0.69 | 0.59 | 0.57 |
| RD | 0.49 | 0.82 | 0.61 | 0.76 | 0.71 | 0.73 | 0.66 | 0.26 | 0.38 | 0.64 | 0.60 | 0.57 | |
| Sonnet | D | 0.49 | 0.85 | 0.63 | 0.89 | 0.62 | 0.73 | 0.66 | 0.38 | 0.48 | 0.68 | 0.62 | 0.61 |
| RD | 0.53 | 0.87 | 0.66 | 0.90 | 0.63 | 0.74 | 0.74 | 0.48 | 0.58 | 0.72 | 0.66 | 0.66 | |
| o3 | RD | 0.75 | 0.66 | 0.70 | 0.67 | 0.86 | 0.76 | 0.73 | 0.61 | 0.67 | 0.72 | 0.71 | 0.71 |
| o4-mini | RD | 0.67 | 0.78 | 0.72 | 0.72 | 0.82 | 0.76 | 0.81 | 0.56 | 0.66 | 0.73 | 0.72 | 0.71 |
| Phi-4 | D | 0.43 | 0.71 | 0.54 | 0.97 | 0.29 | 0.45 | 0.41 | 0.44 | 0.42 | 0.60 | 0.48 | 0.47 |
| RD | 0.35 | 0.86 | 0.50 | 0.56 | 0.05 | 0.09 | 0.65 | 0.27 | 0.38 | 0.52 | 0.39 | 0.32 | |
| Llava-Next | D | 0.39 | 0.90 | 0.55 | 0.68 | 0.47 | 0.55 | 0.00 | 0.00 | 0.00 | 0.36 | 0.46 | 0.37 |
| RD | 0.38 | 0.80 | 0.51 | 0.60 | 0.45 | 0.51 | 0.69 | 0.11 | 0.18 | 0.55 | 0.45 | 0.40 | |
| Llama-3.2 | D | 0.40 | 0.86 | 0.55 | 0.81 | 0.38 | 0.52 | 0.46 | 0.18 | 0.26 | 0.56 | 0.48 | 0.44 |
| RD | 0.38 | 0.93 | 0.54 | 0.67 | 0.12 | 0.20 | 0.53 | 0.18 | 0.27 | 0.52 | 0.41 | 0.34 | |
| Molmo | D | 0.42 | 0.92 | 0.58 | 0.78 | 0.27 | 0.40 | 0.44 | 0.20 | 0.28 | 0.55 | 0.47 | 0.42 |
| RD | 0.38 | 0.80 | 0.52 | 0.58 | 0.28 | 0.38 | 0.36 | 0.16 | 0.22 | 0.44 | 0.41 | 0.37 | |
| InternVL3 | D | 0.66 | 0.76 | 0.70 | 0.88 | 0.74 | 0.80 | 0.65 | 0.66 | 0.66 | 0.73 | 0.72 | 0.72 |
| RD | 0.44 | 0.89 | 0.59 | 0.96 | 0.53 | 0.69 | 0.66 | 0.28 | 0.39 | 0.69 | 0.57 | 0.56 | |
| Qwen2.5 | D | 0.56 | 0.74 | 0.64 | 0.68 | 0.90 | 0.78 | 0.86 | 0.31 | 0.46 | 0.70 | 0.65 | 0.62 |
| RD | 0.41 | 0.91 | 0.57 | 0.88 | 0.43 | 0.58 | 0.81 | 0.24 | 0.37 | 0.70 | 0.53 | 0.51 | |
| DeepSeek | D | 0.62 | 0.36 | 0.45 | 0.50 | 0.76 | 0.60 | 0.43 | 0.39 | 0.41 | 0.52 | 0.50 | 0.49 |
| RD | 0.41 | 0.69 | 0.52 | 0.53 | 0.54 | 0.54 | 0.44 | 0.14 | 0.21 | 0.46 | 0.46 | 0.42 | |
| Support | Neutral | Contradict | Overall | ||||||||||
| P | R | F | P | R | F | P | R | F | P | R | F | ||
| o3 | RD | 0.67 | 0.84 | 0.75 | 0.80 | 0.77 | 0.78 | 0.89 | 0.69 | 0.78 | 0.79 | 0.77 | 0.77 |
| Support | NonSupport | Overall | ||||||||
@@ E.1 Model Performance by Domain
E.2 Few-Shot Experiments
| # Examples | Support | Neutral | Contradict | Overall | |||||||||
| P | R | F | P | R | F | P | R | F | P | R | F | ||
| 4o-mini | 0.40 | 0.90 | 0.56 | 0.66 | 0.42 | 0.51 | 0.74 | 0.10 | 0.18 | 0.6 | 0.47 | 0.42 | |
| 0.41 | 0.90 | 0.56 | 0.81 | 0.12 | 0.22 | 0.73 | 0.48 | 0.58 | 0.65 | 0.50 | 0.45 | ||
| 0.41 | 0.92 | 0.57 | 0.71 | 0.46 | 0.56 | 0.8 | 0.11 | 0.19 | 0.64 | 0.50 | 0.44 | ||
| 0.41 | 0.91 | 0.56 | 0.71 | 0.49 | 0.58 | 0.86 | 0.08 | 0.15 | 0.66 | 0.49 | 0.43 | ||
| GPT-4o | 0.43 | 0.96 | 0.59 | 0.93 | 0.44 | 0.60 | 0.84 | 0.23 | 0.36 | 0.73 | 0.54 | 0.52 | |
| 0.48 | 0.92 | 0.63 | 0.90 | 0.53 | 0.66 | 0.77 | 0.38 | 0.51 | 0.72 | 0.61 | 0.60 | ||
| 0.46 | 0.95 | 0.62 | 0.90 | 0.53 | 0.67 | 0.83 | 0.28 | 0.42 | 0.73 | 0.59 | 0.57 | ||
| 0.45 | 0.95 | 0.62 | 0.90 | 0.51 | 0.65 | 0.86 | 0.28 | 0.42 | 0.74 | 0.58 | 0.56 | ||
| Sonnet | 0.53 | 0.92 | 0.67 | 0.91 | 0.63 | 0.74 | 0.82 | 0.49 | 0.61 | 0.76 | 0.68 | 0.68 | |
| 0.64 | 0.73 | 0.68 | 0.74 | 0.74 | 0.74 | 0.76 | 0.66 | 0.71 | 0.71 | 0.71 | 0.71 | ||
| 0.62 | 0.75 | 0.68 | 0.75 | 0.66 | 0.70 | 0.73 | 0.65 | 0.69 | 0.70 | 0.69 | 0.69 | ||
| 0.60 | 0.83 | 0.69 | 0.83 | 0.69 | 0.76 | 0.77 | 0.60 | 0.67 | 0.73 | 0.71 | 0.71 | ||
We experiment with few-shot prompting (or in-context learning) Brown et al. (2020); Wei et al. (2022) on Biology claims for the decision-only (D) experiments using a subset of closed-source models. Using the methodology described in section 3, we create 45 claims from Cell papers 777These papers are not part of the MuSciClaims evaluation set.. The few-shot examples () are selected randomly from these created claims.
Table 11 presents model performance when prompted to just produce a decision (D). We find that performance improves for all models when they are provided any number of in-context examples. However, the benefits go down as the number of examples goes up.
E.3
MuSciClaims task as a two-class problem
Table 10 presents the results when the main claim verification task of
E.4 Panel Complexity
@@
Appendix F Qualitative Error Analysis
In addition to the diagnostics, we perform qualitative error analysis on a random sample of 100 errors in o3 reasoning (RD) on MuSciClaims. We categorize the errors into the following categories:
1.
Domain Expertise (27%) - Models lack domain expertise, knowledge of related work and common practices of representing data specific to the scientific field.
2.
Visual Understanding (23%) - Models are unable to make the correct inference from the information that they perceive from the figure.
3.
Visual Perception (23%) - Models either miss or pick up data presented in the figure, which can lead to the correct judgment.
4.
Cross Modal Aggregation (17%) - Models incorrectly weight information from one modality over another (such as focus on the caption more than the figure), coming to the wrong conclusion.
5.
Others (10%) - This category contains infrequent error types bucketed together. For instance, models simply misunderstand the caption (textual understanding), or fail to aggregate information from multiple panels (multi-panel aggregation).
Appendix G Prompts Used